# Perturb-and-Project: Differentially Private Similarities and Marginals

### v1

@@ 2 Results
It is known that simply applying the Gaussian mechanism to the input achieves error of the order $O\left(\frac{\log(2/\delta)}{\varepsilon^{2}}\right)\cdot n^{k}$, which is information theoretically optimal when $m\geqslant n^{3k/4}\,,$ up to the dependency on $\varepsilon,\delta\,.$ In the more realistic settings of $m\leqslant n^{3k/4}\,,$ Theorem 2.2 non-trivially improves over this naive algorithm. In comparison, existing polynomial time algorithms (Talwar et al., 2014) could obtain guarantees comparable to Theorem 2.2 only for even $k\,.$ In the odd case, these algorithms yielded an average query wise squared error larger by an $n^{1/2}/(k\log n)$ multiplicative factor. [^2] This gap appears in (Talwar et al., 2014) due to the fact that even order $k$ tensors can be flattened into $k/2\times k/2$ matrices and odd order tensors cannot. Furthermore, the algorithm of (Talwar et al., 2014) required constrained convex optimization methods in which random noise is injected at each iteration.
@@ 5 Guarantees of the perturb-and-project paradigm for differential privacy
It is important to remark that similar results appeared in the risk minimization literature (Chaudhuri & Monteleoni, 2008; Chaudhuri et al., 2011; Jain et al., 2012; Thakurta & Smith, 2013; Duchi et al., 2013; Jain & Thakurta, 2014; Bassily et al., 2014; Ullman, 2015; Chourasia et al., 2021; Nikolov et al., 2013). Specifically, the perturb-and-project framework was studied in (Kifer et al., 2012; Chaudhuri et al., 2011; Dwork et al., 2009), resulting in weaker guarantees than Theorem 5.1. In (Dwork et al., 2015), the special case of Theorem 5.1 for the set of positive semidefinite matrices with bounded nuclear norm was shown. Perhaps, most relevant to our settings is (Talwar et al., 2014). Here the authors studied objective perturbations and bounded in terms of Gaussian complexity of the solution space the change in the objective value for strongly convex functions. Furthermore, (Talwar et al., 2014) also obtained comparable guarantees for the mirror descent algorithm, in which noise is added at each iteration. This difference leads to an important algorithmic consequence: since in Algorithm 1 both the privacy and error guarantees do not depend on the particular steps involved in the computation of $\hat{\bm{\Pi}}$, different techniques may be used to effectively carry out the computation of Algorithm 1 (see Corollary 5.3 and the related discussion).
@@
$\left\lVert A-\Pi(A)\right\rVert^{2}_{2}$ $\geqslant\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}\geqslant$ $\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$ $\leqslant\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$=-2\left\lVert\Pi(A)-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+2\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\leqslant\frac{2}{3}\operatorname*{\mathbb{E}}\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\leqslant\frac{4}{3}\operatorname*{\mathbb{E}}\mathcal{G}(\mathcal{S})\,.$
@@ 5.1 Practical perturb-and-project algorithms
$\operatorname*{\mathbb{E}}\left\lVert\hat{\mathbf{\Pi}}_{t}-\Pi_{\mathcal{S}}(A)\right\rVert^{2}_{2}\leqslant$ $\frac{16\sqrt{\log(2/\delta)}}{3\varepsilon}\cdot\Delta\cdot\left(\mathcal{G}(\mathcal{S})+2^{-t}\left\lVert A\right\rVert^{2}_{2}+n^{-\Omega(1)}\right)\,.$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}$ $\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\leqslant 2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}+2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$ $\leqslant\operatorname*{\mathbb{E}}_{\mathbf{W}}c^{t}\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
@@ Acknowledgments
Tommaso d’Orsi is partially supported by the project MUR FARE2020 PAReCoDi.

### v2

@@ 2 Results
It is known that simply applying the Gaussian mechanism to the input achieves error of the order $O\left(\frac{\log(2/\delta)}{\varepsilon^{2}}\right)\cdot n^{k}$, which is information theoretically optimal when $m\geqslant n^{3k/4}\,,$ up to the dependency on $\varepsilon,\delta\,.$ In the more realistic settings of $m\leqslant n^{3k/4}\,,$ Theorem 2.2 non-trivially improves over this naive algorithm. In comparison, existing polynomial time algorithms (Dwork et al., 2015) could obtain guarantees comparable to Theorem 2.2 only for even $k\,.$ In the odd case, these algorithms yielded an average query wise squared error larger by an $n^{1/2}/(k\log n)$ multiplicative factor. [^2] This gap appears in (Dwork et al., 2015) due to the fact that even order $k$ tensors can be flattened into $k/2\times k/2$ matrices and odd order tensors cannot. Furthermore, the algorithm of (Dwork et al., 2015) required constrained convex optimization methods in which random noise is injected at each iteration.
@@ 5 Guarantees of the perturb-and-project paradigm for differential privacy
It is important to remark that similar results appeared in the risk minimization literature (Chaudhuri & Monteleoni, 2008; Chaudhuri et al., 2011; Jain et al., 2012; Kifer et al., 2012; Thakurta & Smith, 2013; Duchi et al., 2013; Jain & Thakurta, 2014; Bassily et al., 2014; Ullman, 2015; Chourasia et al., 2021; Nikolov et al., 2013). The perturb-and-project framework was first studied in (Nikolov et al., 2013). In (Talwar et al., 2014), the authors studied objective perturbations, bounding the error in terms of Gaussian complexity of the solution space the change in the objective value for strongly convex functions. Furthermore, (Dwork et al., 2015) also obtained comparable guarantees for the Frank-Wolfe algorithm, in which noise is added at each iteration. [^7] Indeed, the argument in (Dwork et al., 2015) is closely related to the proof of Theorem 5.1. This difference leads to an important algorithmic consequence: since in Algorithm 1 both the privacy and error guarantees do not depend on the particular steps involved in the computation of $\hat{\bm{\Pi}}$, different techniques may be used to effectively carry out the computation of Algorithm 1 (see Corollary 5.3 and the related discussion).
@@
$\left\lVert A-\Pi(A)\right\rVert^{2}_{2}$
$\geqslant\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle$
$+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$\geqslant\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2}$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$+\frac{1}{2}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant\langle\nabla\left\lVert A-\Pi(A+\mathbf{W})\right\rVert^{2}_{2},\Pi(A)-\Pi(A+\mathbf{W})\rangle$
$+\langle\nabla\left\lVert A+\mathbf{W}-\Pi(A)\right\rVert^{2}_{2},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$=-2\left\lVert\Pi(A)-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+2\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle\,.$
@@
$\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant\frac{2}{3}\operatorname*{\mathbb{E}}\langle\mathbf{W},\Pi(A+\mathbf{W})-\Pi(A)\rangle$
$\leqslant\frac{4}{3}\operatorname*{\mathbb{E}}\mathcal{G}(\mathcal{S})\,.$
@@ 5.1 Practical perturb-and-project algorithms
$\operatorname*{\mathbb{E}}\left\lVert\hat{\mathbf{\Pi}}_{t}-\Pi_{\mathcal{S}}(A)\right\rVert^{2}_{2}\leqslant$ $\frac{16\sqrt{\log(2/\delta)}}{3\varepsilon}\cdot\Delta$
$\cdot\left(\mathcal{G}(\mathcal{S})+2^{-t}\left\lVert A\right\rVert^{2}_{2}+n^{-\Omega(1)}\right)\,.$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}$ $\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
$\leqslant 2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$+2\operatorname*{\mathbb{E}}_{\mathbf{W}}\left\lVert\Pi(A+\mathbf{W})-\Pi(A)\right\rVert^{2}_{2}$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{W}}$ $\left\lVert\Pi_{t}(A+\mathbf{W})-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
$\leqslant\operatorname*{\mathbb{E}}_{\mathbf{W}}c^{t}\left\lVert A+\mathbf{W}-\Pi(A+\mathbf{W})\right\rVert^{2}_{2}$
@@ Acknowledgments
The authors thank A. Nikolov for useful discussions concerning $k$-way marginal queries and the perturb-and-project framework. Tommaso d’Orsi is partially supported by the project MUR FARE2020 PAReCoDi.
Impact statement
This paper presents work whose goal is to advance the field of differentially private machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.